OpenStudy (xoomd):

What is the equation of the line that passes through the points (−2, 3) and (2, 7)?

OpenStudy (3mar):

may I help?

OpenStudy (xoomd):

please

OpenStudy (3mar):

ok

OpenStudy (3mar):

This is the equation form of the line passes through two points: \[\frac{ y_2-y_2 }{ x_2-x_1 }=\frac{ y=y_1 }{ x-x_1 }\] All you need is to substitute with the two points in the equation

OpenStudy (xoomd):

So I found the slope.. and it was 1 Do i need that to plug in?

OpenStudy (3mar):

yes the slope is 1 correct

OpenStudy (xoomd):

im confused about how to use the formula you gave

OpenStudy (3mar):

one minute please

OpenStudy (xoomd):

okay

OpenStudy (3mar):

yes i am here

OpenStudy (3mar):

I am sorry for late!

OpenStudy (xoomd):

no problem

OpenStudy (3mar):

the left side is for slope the right side is to y and x

OpenStudy (3mar):

as the slope is 1, put it instead of the left side.

OpenStudy (xoomd):

I wrote an equation that looked like this : Y-3=1(x-(-2)

OpenStudy (3mar):

Great! You did it. Some arrangement

OpenStudy (xoomd):

but then I wound up getting y+1x-1

OpenStudy (xoomd):

Y=1x-1 I mean

OpenStudy (3mar):

\[\frac{ y-3 }{ x-(-2) }=1\] \[y-3=x+2\] \[y=x+5\] and this this the graph of that line

OpenStudy (xoomd):

The possible answers i was given were: x − y = −1 x − y = −2 x − y = − 5 x − y = − 6

OpenStudy (3mar):

got it? If you stuck to anything, tell me.

OpenStudy (xoomd):

Out of these choices, is it the first one? x − y = −1 x − y = −2 x − y = − 5 x − y = − 6

OpenStudy (3mar):

Easy if we rearrange the last form, we get \[y=x+5\] \[y-y=x+5-y\] \[x-y-5=0\] \[x-y-5+5=0+5\] \[x-y=-5\] Hope that helps

OpenStudy (xoomd):

Ohh I see, Thank you!

OpenStudy (3mar):

Thank you for learning!!

OpenStudy (3mar):

and Thank you for the medal!

OpenStudy (sshayer):

y-y1=m(x-x1) \[m=\frac{ y2-y1 }{ x2-x1 }\]